

********************************************************************
UARS WEEKLY STATUS REPORT - 26 NOV 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
Happy Thanksgiving!
***********
SUSIM
***********
Until the newest telemetry outages, SUSIM successfully gathered its
normal solar data this past week without apparent problems.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
Investigation of the tape recorder anomaly has continued this
week and the MPG has been involved with this activity. MPG
activity has been concentrated on evaluation of the potential
impact on instrument operations and science data collection.
***********
CDHF
***********
The CDHF continues to notify all PI Software Representatives when Data Gaps
occur in the LEVEL0 data due to the UARS Spacecraft Anomaly.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
9 days CLA (7.0) PROD (L1-L3) (476R-400C-19F=57L)
10 days GPI (4.05) PROD (CLAES L3B) (334R-317C-2F=15L)
16 days WIN (5.11A) PROD (L1-L3) (520R-19C-2F=499L)
DAAC DATA:
Data created in the month of September 1999 has been previewed for HALOE,
PEM, NMC, and UKMO.
UARS CDHF REHOST:
Testing SUSIM V7.1 on the Rehost System.
Preparing to test SOLSTICE V4.62 on the Rehost System.
***********
FOT
***********
REPORTING PERIOD
This report covers 13 November, 1999 (Orbit 44660, GMT Day 317) through 19
November 1999 (Orbit 44764, GMT Day 323).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 317/2327 44675 AXIS-1 Off
WINDII 319/0026 44686 Off
The beta angle ranged between the angles of 60.8 degrees to 74.9 degrees for
this report period. The beta angle is now decreasing toward a minimum angle of
-2.0 degrees on 08 December 1999 (DOY 342).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 0.3 to 1.0 degrees C. Additional battery performance data for each
battery on 12 November 1999 and 19 November 1999 are provided in App D.
The clock error ranged from 10.8 to -9.9 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 13
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
% HRDI = 08
% FOT = 05
% PEM = 01
% WINDII = 01
% SUSIM = 01
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
% FOT = 03
Three (3) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
% Transponder A: 99/284 - 2287.496981
% Transponder B: 99/265 - 2287.498700
GROUND SYSTEM OPERATION
Two (2) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 99/322 show the observatory data loss to be 150 hours, 56
minutes, 50.105 seconds (an increase of 15 minutes 31.008 seconds since the
last report period). This is a 0.2112 percent data loss which equals a 99.7888
percent data capture for the mission.
The increase of 15 minutes 31.008 seconds occurred over seven (7) days:
7 02 minutes 15.616 seconds on Day 316 due to a S/C TR problem.
7 01 minutes 43.712 seconds on Day 317 due to a S/C TR problem.
7 00 minutes 26.496 seconds on Day 318 due to a S/C TR problem.
7 00 minutes 19.008 seconds on Day 319 due to a S/C TR problem.
7 03 minutes 03.872 seconds on Day 320 due to a S/C TR problem.
7 01 minutes 44.896 seconds on Day 321 due to a S/C TR problem.
7 05 minutes 57.408 seconds on Day 322 due to a S/C TR problem.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The Leonid operation was successful. This event occurred on Wed, Nov 17, 1999
DOY 321. The Flight Operations Team disabled the Fixed Head Star Tracker
shutter reopen TMON's during the storm period. In addition, no supports were
taken during the storm. All science instruments that wanted to be on during
the storm remained on. The TMON's were reconfigured after the storm. All in
all the event went well with no incidents to report.
As of Fri., Nov.19, 1999, DOY 323, the FOT began scheduling all future TDRS
contacts according to what will be considered the new "normal TDRS scheduling
scenario". Under this scenario, the FOT will attempt to schedule one 20-minute
SSA event every other orbit. These 20-minute events will be used for NBTR
playbacks. In between orbits will be filled with either 15-minute SSA events,
or 15-minute MA events. This equates to one event (either SSA or MA) every
satellite orbit. There will be minor deviations to this guideline, depending
on TDRS availability. The focus of this scenario is to eliminate frequent
"short" NBTR playbacks (i.e. 7 minutes or less), thus utilizing a greater
portion of the available record cycle of the NBTR. As usual, the FOT will
attempt to schedule all TDRS events during spacecraft daylight periods,
whenever possible.
YEAR 2000 OPERABILITY SUPPORT
All areas identified to or by the Flight Operations Team for Y2K testing have
been completed. There are no other areas identified at this time that have not
been Y2K certified.
MEETINGS
DOY 319, Mon. 11:15 am, Nov 15, 1999 - Teleconference with Tim Ohrman with the
company, L3 that now produces the NASA Standard Tape Recorder. He said that
when NBTR-A failed that it was totally failed in the TU, and was completely
unusable. He said that it wouldn't hurt to turn it on, but we would most
likely not get anything out of it. Described to Tim the test plans that the
Flight Operations Team (FOT) put together. His view were in concurrence with
our own. The only capability that he felt was not available was the Erase
function while in wind mode. We agreed to send him some plots for him and one
of his engineers to review.
DOY 319, Mon. 1:00 PM, Nov 15, 1999 - Meeting held to discuss the status of
NBTR B. The Flight Operations Plan was presented to all participants. The FOT
recommended that 4 Mux Mode be continue for a 10 period. After that period
(Nov 9 -19) real-time scheduling would return to a normal. The FOT received
approval asked to continue to pursue the anomaly investigation and continue
with any tests necessary.
Additional contracted support attended this meeting. Data was presented that
supported the functionality of an Erase capability in the wind mode along with
the hex command. The FOT compared the RIU, Channel and Hex code of UARS
commands to the provided commands from Landsat 4/5 and they matched. This
generates confidence that the hex command that we have is correct provided the
function exists. As mentioned previously, the erase command is essential to
PASS #2 operations and the complete capture of science data.
FUTURE SIGNIFICANT EVENTS
No specific date was agreed upon for UARS Safehold configuration discussions.
Instead, the Supporting Engineers involved will be contacted every two weeks to
determine if either the Flight Operations Team or Supporting Engineers have any
new developments that should be discussed in a group. It is expected that one
more meeting will be held before Y2k as an update to UARS safehold expectations
for the new century.
UARS has tested and prepared for the Year 2000 event. Special considerations
have been made in preparation for a non nominal 1999 to 2000 transition.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 19 November 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
44675 317/2327 99-373 PEM AXIS 1 Power OFF
44686 318/1737 99-374 OBC Clk Adj fr 20.6 to 41.2 msec/day
44690 319/0026 99-375 WINDII Power OFF
44694 319/0705 99-376 HRDI SAFE MODE
44704 319/2150 99-377 HRDI SAFE MODE
44717 320/1913 99-378 HRDI HRDI RESTART
44723 321/0540 99-379 HRDI HRDI to SAFEHOLD - Hi Temp
44733 321/2014 99-380 SUSIM MONITOR RATE
44734 321/2205 99-381 OBC Clk Adj 41.2 to 20.6 msec/day
44734 321/2205 99-382 FHST-2 TMON 37 DISABLED
44739 322/0636 99-382 FHST-2 TMON 37 ENABLED
44740 322/0857 99-383 HRDI HRDI to Safehold - Hi Temps
44753 323/0458 99-384 HRDI Power OFF(low pwr)
44755 323/0818 99-385 OBC Clk Adj fr 20.6 to 41.2 msec/day
44756 323/0904 99-386 HRDI HRDI to Safehold position
44757 323/1041 99-387 HRDI HRDI to Safehold pos.- functional
___________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 19 November 1999
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
QTZEPSIF1 Y-LO ZEPS INTERFACE TEMP 1 44683
HRPTSBARTMP Y-LO HRDI TELESCOPE BARREL TEMP 44692-95
HRPTSBARTMP R-LO HRDI TELESCOPE BARREL TEMP 44692-44704
HRPTSBARTMP Y-LO HRDI TELESCOPE BARREL TEMP "44705-09,18"
HRPTSBARTMP R-LO HRDI TELESCOPE BARREL TEMP 44709-17
HRPTSBARTMP Y-LO HRDI TELESCOPE BARREL TEMP 44725
HRPTSBARTMP R-LO HRDI TELESCOPE BARREL TEMP 44720-25
NBTUTTEMP2 Y-LO 38.O DEGREES "44728,9"
HRPTSBARTMP Y-LO HRDI TELESCOPE BARREL TEMP "44738-40,43,46,48"
HRPTSBARTMP R-LO HRDI TELESCOPE BARREL TEMP 44740
HRPTSBARTMP R-LO HRDI TELESCOPE BARREL TEMP 44752-58
NBTUTTEMP2 Y-LO NBTR TU TEMP 44750
HRPTSBARTMP Y-LO HRDI TELESCOPE BARREL TEMP 44759
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
No Alerts to report this report period.
_________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 19 November 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
44686 318/17:45 (14 Nov 99) from 20.6 to 41.2 msec/day
44734 321/22:05 (17 Nov 99) from 41.2 to 20.6 msec/day
44755 323/08:25 (19 Nov 99) from 20.6 to 41.2 msec/day
__________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 19 November 1999
12 Nov 1999 - GMT Day 316 Beta = 60.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- -------- --------- ---- ------- ------
1 -89.6/-106.4 -3.05/-4.94 +0.0/+0.0 16.3 0.000 0.0
2 +28.0/-11.2 +5.47/+3.87 +29.6/-15.2 25.3 1.086 21.2
3 +33.6/-22.4 +4.51/+2.60 +29.2/-15.2 25.1 1.083 21.0
19 Nov 1999 - GMT Day 323 Beta = 74.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- ------- --------- ----- ------ ------
1 -67.2/-89.6 -2.42/-3.36 +0.0/+0.0 16.3 0.000 0.0
2 +22.4/+0.0 +3.23/+1.66 +16.8/-14.8 27.5 1.037 13.4
3 +22.4/-22.4 +2.29/+1.34 +16.0/-14.4 27.4 1.035 12.9
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 19 November 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-081 FOT 44670 317/1639 14170 Generic Late Acquisition
99-083 FOT 44746 322/1837 14171 Generic Drop Out on Return Link
99-081 FOT 44749 322/2204 14170 Generic Late Acquisition
99-081 FOT 44752 323/0353 14170 Generic Late Acquisition
99-084 FOT 44757 323/1041 N/A HRDI BARREL TEMP 56.08
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-001 FOT 44670 317/1639 14170 Generic Late Acquisition
99-083 FOT 44746 322/1837 14171 Generic Drop Out on Return Link
99-081 FOT 44749 322/2204 14170 Generic Late Acquisition
99-081 FOT 44752 323/0353 14170 Generic Late Acquisition
99-084 FOT 44757 323/1041 N/A HRDI BARREL TEMP 56.08
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
No new reports generated
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-064 FOT 44135 10/8-9/99 N/A HRDI Cal Lamp Errors
99-071 FOT 44398 299/1212 21512 TPF Not Receiving 512 PB
99-081 FOT November 14170 Generic Late Acquisition
SEVNTFW99/47